OpenStudy (anonymous):

Seven girls on a bus are being driven to school. Each girl has seven bags. For each bag there are 7 big cats and for every big cat there are 7 little cats. How many legs are on the bus?

OpenStudy (anonymous):

geerky42 has a small vag

geerky42 (geerky42):

wow how did you know? I wondering...

OpenStudy (anonymous):

i licked it once

OpenStudy (anonymous):

So there are 7 girls on the bus, they have 7*2=14 legs They all have seven bags, thus 7*7=49 bags total. (bags dont have legs) There are 7 big cats for each bag: 7*7*7= 343 they all have 4 legs 343*4=1372 legs Little cats 7*7*7*7=2401 with 4 legs each 9604 Add all the above: 14+1372+9604=10990 legs Bonus question: How do you fit so many things into 7 buses. Or is it a trick question and the answer is 14? Because the cats are not on the buses :)

OpenStudy (anonymous):

@andras Forgot to count the driver. :p